A nickel coil has a resistance of 13 Ω at 50°C. If the temperature coefficient of resistance at 0°C is 0.006/°C, the resistance at 0°C is


1
Expert's answer
2021-08-23T12:40:01-0400

The resistance at 50°:


"R=R_0(1+kt),\\
R_0=\frac{R}{1+kt}=10\Omega."